OpenStudy (anonymous):

Do you need to know all of the factors of a number in order to determine if it is prime or composite? @MATH

OpenStudy (jamesj):

No, if you find one factor (that isn't 1 or the number itself), that's sufficient. For example n = 123473456742357832458902 This number has a factor of 2, and now we can immediately say that n is composite. We don't need to find all of the other factors.

OpenStudy (anonymous):

what abt a prime number?

OpenStudy (jamesj):

Testing whether a number is prime is a pain. The simplest way to test whether a number n is prime is to check whether every integer up to \( \sqrt{n} \) is a factor. For example, suppose n = 1001. \( \sqrt{n} = 31.6385.. \) so we will need to check all the primes from 2 to 31.

OpenStudy (jamesj):

For example, does 2 divide 1001? We write 2 | 1001. No, because if it did, the last digit would be 0,2,4,6 or 8. Next. 3 | 1001? No, because if it did, the sum of the digits of 1001 would be divisible by 3. Next 5 | 1001 ? No, because if it did, the last digit would be 5 or 0 Next 7 | 1001 ? Yes. 1001 = 7 * 143. Hence 1001 is not prime.

OpenStudy (jamesj):

If however we had kept on going and none of the primes between 2 and 31 divided n = 1001, then we could conclude 1001 was prime

OpenStudy (anonymous):

LOL HEHE

OpenStudy (jamesj):

lol. Ah, trust me. I used to lecture mathematics at university.

OpenStudy (jamesj):

No. Read it again and internalize it.

OpenStudy (jamesj):

In my head, the explanation is about 1/2 sec long. If you fully understand it, it'll be that long for you too. I wrote it the way I did so you could understand it. I hope you do. Making it shorter will make it harder to understand, not easier.

OpenStudy (jamesj):

You understand it and rewrite it to whatever size you need. I've got to move on.